Given = î + ĵ − 2 and = î − 4 ĵ − , calculate...

Given M with arrow = î + ĵ − 2 k and N with arrow = î − 4 ĵk, calculate the vector product M with arrowN with arrow.
--i+---j+---k
Answer #1

M = i +9 - 2k, = î– 45 – k MXN = 1 -2 1 -4 -* | = i (1X(-1)-(-2) (-4)] - ĵ (C) (-)-(-2) CU] t² Clic-4)-(1303 M N = î (-1-8) –

=-9i+(-1)j+(-5)k